# Semi-Infinite Programs for Robust Control and Optimization: Efficient Solutions and Extensions to Existence Constraints

### v1

@@ 2.1 Obstacle Avoidance
For (5) to guarantee that the constraint of (3) is satisfied, it is simply necessary and sufficient for there to exist a choice of $\mbf{s}_{k}$ satisfying (5) and (2.1) for every possible realization of the uncertainty. However, it is trivial to observe that while (5) and (2.1) can admit non-unique solutions $\mbf{s}_{k}$, it is only necessary to prove the existence of a single solution for the original constraint of (3) to be satisfied. This is a problem that cannot be solved using existing local reduction-based SIP solvers.
@@ 2.2 Input Saturation
The same procedure can then be applied to the two other conditions in (11) to recover (1). Since every step between (10) and (2.2) is a two-way implication, (8) and (1) must be equivalent, completing the proof. ∎
@@ 3.2 Robust Optimal Control Problem
Using Assumption 2, (3.2) can be rewritten as
@@
where $\gamma\in\Gamma\subset\mathbb{R}$ is introduced as an upper bound on the objective value. Djelassi and Mitsos (2021) introduces a method to deal with the existence constraint of (30b) explicitly, but relies on global optimization techniques and is unsuitable for adaptation to local reduction. Instead, we propose rewriting (3.2) as
@@
From Assumptions 3 and 5, we know that $(\theta,w,\gamma)\mapsto f(\theta,w,z_{p}^{*}(\theta,w))-\gamma$ is continuous. Similarly, using Assumption 4 for the compactness of $\mathcal{S}$, it follows from Clarke (1975) that $(\theta,w)\mapsto\min_{s\in\mathcal{S}}g(\theta,w,z_{p}^{*}(\theta,w),s)$ is also continuous. Finally, given that $\Theta$ and $\mathcal{W}$ are compact by definition and $\Gamma$ is compact under Assumption 6, the problem of (3.2) can be rewritten as
@@ 4 Local Reduction Algorithm
While the problem of (3.2) is useful in proving that local reduction can be employed to solve the desired min-max-min problem, it is far from ideal for use with numerical solvers. Notably, the function $z_{p}^{*}$ is difficult to find explicitly, rendering the exact representation of (3.2) close to impossible. Instead, we apply the local reduction algorithm described in Zagorowska et al. (2024) optimizing for the tuple $(\theta,\gamma)$ under a finite number of scenarios $w_{i}\in\mathbb{W}_{j}$ where $\text{card}\left(\mathbb{W}_{j}\right)=j$. Under any given policy $(\theta,\gamma)$, the worst-case uncertainty $w$ for the constraint of (31c) is found through
@@
which is a nested SIP that can be solved directly through local reduction under Assumption 7, since (4) does not have any inner optimization problems inside its constraints in the way that (3.2) does.
@@
$g(\theta_{i},w_{i},z_{p,i},s_{i})$ $\leq 0$ (36b)
$f(\theta_{i},w_{i},z_{p,i})-\gamma$ $\leq 0$ (36c)
$d(\theta_{i},w_{i},z_{p,i},z_{m,i})$ $=0$ (36d)
$e(\theta_{i},w_{i},z_{p,i},z_{m,i})$ $\leq 0$ (36e)

### v2

@@ 2.1 Obstacle Avoidance
For (5) to guarantee that the constraint of (3) is satisfied, it is simply necessary and sufficient for there to exist a choice of $\mbf{s}_{k}$ satisfying (5) and (6) for every possible realization of the uncertainty. However, it is trivial to observe that while (5) and (6) can admit non-unique solutions $\mbf{s}_{k}$, it is only necessary to prove the existence of a single solution for the original constraint of (3) to be satisfied. This is a problem that cannot be solved using existing local reduction-based SIP solvers.
@@ 2.2 Input Saturation
The same procedure can then be applied to the two other conditions in (11) to recover (9). Since every step between (10) and (14) is a two-way implication, (8) and (9) must be equivalent, completing the proof. ∎
@@ 3.2 Robust Optimal Control Problem
Using Assumption 2, (27) can be rewritten as
@@
where $\gamma\in\Gamma\subset\mathbb{R}$ is introduced as an upper bound on the objective value. Djelassi and Mitsos (2021) introduces a method to deal with the existence constraint of (30b) explicitly, but relies on global optimization techniques and is unsuitable for adaptation to local reduction. Instead, we propose rewriting (30) as
@@
From Assumptions 3 and 5, we know that $(\theta,w,\gamma)\mapsto f(\theta,w,z_{p}^{*}(\theta,w))-\gamma$ is continuous. Similarly, using Assumption 4 for the compactness of $\mathcal{S}$, it follows from Clarke (1975) that $(\theta,w)\mapsto\min_{s\in\mathcal{S}}g(\theta,w,z_{p}^{*}(\theta,w),s)$ is also continuous. Finally, given that $\Theta$ and $\mathcal{W}$ are compact by definition and $\Gamma$ is compact under Assumption 6, the problem of (31) can be rewritten as
@@ 4 Local Reduction Algorithm
While the problem of (31) is useful in proving that local reduction can be employed to solve the desired min-max-min problem, it is far from ideal for use with numerical solvers. Notably, the function $z_{p}^{*}$ is difficult to find explicitly, rendering the exact representation of (31) close to impossible. Instead, we apply the local reduction algorithm described in Zagorowska et al. (2024) optimizing for the tuple $(\theta,\gamma)$ under a finite number of scenarios $w_{i}\in\mathbb{W}_{j}$ where $\text{card}\left(\mathbb{W}_{j}\right)=j$. Under any given policy $(\theta,\gamma)$, the worst-case uncertainty $w$ for the constraint of (31c) is found through
@@
which is a nested SIP that can be solved directly through local reduction under Assumption 7, since (35) does not have any inner optimization problems inside its constraints in the way that (31) does.
@@
$g(\theta,w_{i},z_{p,i},s_{i})$ $\leq 0$ (36b)
$f(\theta,w_{i},z_{p,i})-\gamma$ $\leq 0$ (36c)
$d(\theta,w_{i},z_{p,i},z_{m,i})$ $=0$ (36d)
$e(\theta,w_{i},z_{p,i},z_{m,i})$ $\leq 0$ (36e)